# Iterated Denoising Energy Matching for Sampling from Boltzmann Densities

### v1

@@ Abstract
Efficiently generating statistically independent samples from an unnormalized probability distribution, such as equilibrium samples of many-body systems, is a foundational problem in science. In this paper, we propose Iterated Denoising Energy Matching (iDEM), an iterative algorithm that uses a novel stochastic score matching objective leveraging solely the energy function and its gradient—and no data samples—to train a diffusion-based sampler. Specifically, iDEM alternates between (I) sampling regions of high model density from a diffusion-based sampler and (II) using these samples in our stochastic matching objective to further improve the sampler. iDEM is scalable to high dimensions as the inner matching objective, is simulation-free, and requires no MCMC samples. Moreover, by leveraging the fast mode mixing behavior of diffusion, iDEM smooths out the energy landscape enabling efficient exploration and learning of an amortized sampler. We evaluate iDEM on a suite of tasks ranging from standard synthetic energy functions to invariant $n$-body particle systems. We show that the proposed approach achieves state-of-the-art performance on all metrics and trains 25×2-5$\times$2 - 5 × faster, which allows it to be the first method to train using energy on the challenging $55$-particle Lennard-Jones system [^1] Code for iDEM is available at https://github.com/jarridrb/dem..
@@ 1 Introduction
[…]
@@ 3.1 Denoising diffusion with a Boltzmann target (C1)
$\left|\frac{1}{K}\sum_{i}\exp(-{\mathcal{E}}(x_{0\mid t}^{(i)}))-\exp(-{\mathcal{E}}_{t}(x_{t}))\right|\leq C\sqrt{\frac{\log(\frac{2}{\delta})}{K}}\;\text{and}\;\left\|\frac{1}{K}\sum_{i}\nabla\exp(-{\mathcal{E}}(x_{0\mid t}^{(i)}))-\nabla\exp(-{\mathcal{E}}_{t}(x_{t}))\right\|\leq C\sqrt{\frac{\log(\frac{2}{\delta})}{K}}$ (10)
@@
[…]
@@ 3.3 Incorporating symmetries in iDEM
Note that in the first line we used that $x_{0\mid t}^{(i)}\sim\overline{{\mathcal{N}}}(x_{t},\sigma^{2}_{t})$ is equivalent to $g\circ x_{0\mid t}^{(i)}\sim\overline{{\mathcal{N}}}(g\circ x_{t},\sigma^{2}_{t})$. In practice, we can easily implement an equivariant ${\mathcal{S}}_{K}(x_{t},t)$ by replacing the standard normal distribution with a normal distribution that has zero center of mass. Note that a standard normal is already rotation and permutation invariant and an orthogonal action induces no change in volume as the determinant is $1$. Whilst it is not possible to define a translation invariant measure on ${\mathbb{R}}^{d}$ we can still achieve this symmetry by constructing a normal distribution that has zero center of mass. Intuitively, this is a projection of the density in ${\mathbb{R}}^{d}$ to ${\mathbb{R}}^{d-1}$ and this subspace is translation invariant (Köhler et al., 2020; Garcia Satorras et al., 2021; Midgley et al., 2023a). Finally, to use symmetries within our iDEM algorithm we also need to take the diffusion sampler $s_{\theta}$ to be equivariant to the product group. While this design choice is necessary it does not come with any loss of generality as there always exists an equivariant map between two group invariant distributions on ${\mathbb{R}}^{d}$ (Bose et al., 2021).
@@ 4 Experimental results
We evaluate iDEM on multiple unnormalized densities including synthetic and $\mathrm{SE(3)}\times\mathbb{S}_{n}$-equivariant $n$-body particle systems of varying complexity as examples of scientific applications. For our metrics, we report both sample-based metrics, such as 2-Wasserstein and Total Variation (TV) distance that assess mode coverage, as well as the standard negative log-likelihood (NLL).
@@ 4.1 Main results
Table 2: Sampler performance with mean $\pm$ standard deviation over 3 seeds for negative log-likelihood (NLL), Total Variation (TV), and 2-Wasserstein metrics ($\mathcal{W}_{2}$). *** indicates divergent training. Bold via Welch’s two sample t-test $p<0.1$. See §F.2 for more details.
@@
Algorithm $\downarrow$ NLL TV ${\mathcal{W}}_{2}$ NLL TV ${\mathcal{W}}_{2}$ NLL TV ${\mathcal{W}}_{2}$ NLL TV ${\mathcal{W}}_{2}$
FAB (Midgley et al., 2023b) 7.14±$\pm 0.01$ 0.88±$\pm 0.02$ 12.0±$\pm 5.73$ 7.16±$\pm 0.01$ 0.09±$\pm 0.00$ 2.15±$\pm 0.02$ 17.52±$\pm 0.17$ 0.04±$\pm 0.00$ 4.35±$\pm 0.01$ 200.32±$\pm 62.3$ 0.24±$\pm 0.09$ 18.03±$\pm 1.21$
PIS (Zhang & Chen, 2022) 7.72±$\pm 0.03$ 0.92±$\pm 0.01$ 7.64±$\pm 0.92$ 7.19±$\pm 0.01$ 0.09±$\pm 0.00$ 2.13±$\pm 0.02$ 47.05±$\pm 12.46$ 0.25±$\pm 0.01$ 4.67±$\pm 0.11$** *** ***
DDS (Vargas et al., 2023) 7.43±$\pm 0.46$ 0.82±$\pm 0.02$ 9.31±$\pm 0.82$ 11.27±$\pm 1.24$ 0.16±$\pm 0.01$ 2.15±$\pm 0.04$** *** *** *** *** ***
pDEM (ours) 7.10±$\pm 0.02$ 0.82±$\pm 0.02$ 12.20±$\pm 0.14$ 7.44±$\pm 0.05$ 0.13±$\pm 0.00$ 2.11±$\pm 0.03$ 18.80±$\pm 0.48$ 0.06±$\pm 0.02$ 4.21±$\pm 0.06$** *** ***
iDEM (ours) 6.96±$\pm 0.07$ 0.82±$\pm 0.01$ 7.42±$\pm 3.44$ 7.17±$\pm 0.00$ $\pm 0.01$ 2.13±$\pm 0.04$ 17.68±$\pm 0.14$ 0.04±$\pm 0.01$ 4.26±$\pm 0.03$ 125.86±$\pm 18.03$ 0.09±$\pm 0.01$ 16.128±$\pm 0.071$
@@
Table 3: Training time results in hours excluding evaluation time. *** denotes divergent training runs.
@@
PIS (Zhang & Chen, 2022) 4.11 11.29 17.36 ***
DDS (Vargas et al., 2023) 1.81 5.65 *** ***
pDEM (ours) 0.36 1.40 1.79 ***
@@ 6 Conclusion
J.R., N.M., and A.T. initially conceived the idea of a stochastic off-policy continuous regression objective. A.J.B., T.A., G.G., and A.T. independently approached the problem through the lens of active inference with a generative model in a bi-level iterative scheme. Experiments were primarily led by T.A., J.R., A.T. (iDEM) and S.M., P.L., C.H.L., M.S. (ablations and baselines). G.G., A.J.B., and N.M. led the development of the theory. S.R., G.G., and Y.B. guided the project. A.J.B. and N.M. drove the writing of the paper, with contributions from all other authors. All authors contributed to designing the experiments.
@@ D.5 Boltzmann generators
A Boltzmann generator (BG) samples from $\mu_{\text{target}}$ by combining an exact-likelihood model $q_{\theta}$, typically a Normalizing Flow, and an algorithm to reweight model generated samples using $\mu_{\text{target}}$. During training a BG’s flow is learned by minimizing a convex combination of both the forward and reverse KL-divergence. Thus, to compute the full training loss we need a dataset of ground truth samples from $\mu_{\text{target}}$ as it is required under the forward KL. As a result, BG’s are ill-suited in settings with limited or no data as training only with the reverse KL is prone to mode-seeking behavior. Given a trained flow $q_{\theta}$ we can reweight observables $f(x)$ to be under $\mu_{\text{target}}$ by using the IS estimator.
@@ F.1 Experimental Setup
F.2 Table 2
Tab. 2 depicts the main experimental sample quality results for various methods on all datasets. In this section, we explain the methodology for each experiment. We note that it was difficult to train PIS and DDS for the high-dimensional Lennard-Jones tasks. 2/3 seeds of DDS diverged for LJ-13.
@@ F.2 Table 2
Total Variation. The total variation metric on the Gaussian mixture model dataset is taken over 200 bins in each dimension (${}^{2}$) total bins. This is possible in low dimensions but does not scale well to high dimensions, requiring an exponential number of bins. Therefore for the larger equivariant datasets, we take the total variation distance over the distribution of the interatomic distances of the particles, see e.g., Fig. 4, which scales well with dimension and shows how well the distribution of energies matches that of the test data.
2-Wasserstein distance $\mathcal{W}_{2}$. We also use the standard 2-Wasserstein distance between empirical samples from the sampler and from the ground truth dataset. The 2-Wasserstein distance is defined as
@@
where $\pi$ is the transport plan with marginals constrained to $\mu$ and $\nu$ respectively. In practice, we use the Hungarian algorithm as implemented in the Python optimal transport package (POT) (Flamary et al., 2021) to solve this optimization for discrete samples. We use the Euclidean ground distance $d(x,y)=\|x-y\|_{2}$ for simplicity.
@@ F.4.1 Gaussian Mixture Model
We use use a test set of size 1000 sampled with torch.random.seed(0) following prior work.
@@ G.1 Additional metrics
We report here additional quantitative metrics to supplement the main results presented in §4.1. In Tab. 5 we report the NLL values that are computed between the original method and the CFM model that is trained using the samples of each method.
@@ G.3 Further ablations
$=-\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{x_{0|t}\sim\mathcal{N}(x_{t},\sigma_{t}^{2})}\left[\mathcal{E}(x_{0|t})\|x_{0|t}-x_{t}\|\right],$
@@
We investigate the behavior of each estimate in Figure 10, examining the MSE between the score estimated by each technique and the true score on the GMM task as a function of the number of MC samples. Unfortunately, we observe that despite the Jensen estimate’s attractiveness the score estimates it yields have a large MSE which does not readily reduce as the number of MC samples increases indicating a significant bias. On the other hand, the unbiased estimate suffers from extremely large variance due to it having to work directly with the exponentiated energies $\exp(-\mathcal{E}(x_{t}))$. This results in every estimate before 500 MC samples resulting in exclusively NaN or infinite estimates, rendering the estimate unusable in practice. The only estimate which is well-behaved and approaches the true score is the LogSumExp estimate, whose use we advocate throughout this work.
@@
In Fig. 11 (left) we study the log-bias as a function of ${\mathcal{E}}$, by taking a point from a GMM mode and linearly following a direction away from the data itself which corresponds to $\exp(-{\mathcal{E}}(x))\approx 0$. As observed, we notice we need a few MC samples in high data regions. Additionally, the log-bias increases further from data for a fixed $K$ and drops when we increase $K$. This result empirically confirms §3.1 where we need more MC samples to have a proper estimation of the score.

### v2

@@ Abstract
Efficiently generating statistically independent samples from an unnormalized probability distribution, such as equilibrium samples of many-body systems, is a foundational problem in science. In this paper, we propose Iterated Denoising Energy Matching (iDEM), an iterative algorithm that uses a novel stochastic score matching objective leveraging solely the energy function and its gradient—and no data samples—to train a diffusion-based sampler. Specifically, iDEM alternates between (I) sampling regions of high model density from a diffusion-based sampler and (II) using these samples in our stochastic matching objective to further improve the sampler. iDEM is scalable to high dimensions as the inner matching objective, is simulation-free, and requires no MCMC samples. Moreover, by leveraging the fast mode mixing behavior of diffusion, iDEM smooths out the energy landscape enabling efficient exploration and learning of an amortized sampler. We evaluate iDEM on a suite of tasks ranging from standard synthetic energy functions to invariant $n$-body particle systems. We show that the proposed approach achieves state-of-the-art performance on all metrics and trains 25×2-5$\times$2 - 5 × faster, which allows it to be the first method to train using energy on the challenging $55$-particle Lennard-Jones system.
@@ 1 Introduction
Can we find a scalable sampler that can learn from $\mathcal{E}(x)$ and $\nabla{\mathcal{E}}$ while achieving high mode-coverage of $\mu_{\text{target}}$?
@@ 3.1 Denoising diffusion with a Boltzmann target (C1)
$\left|\frac{1}{K}\sum_{i}\exp(-{\mathcal{E}}(x_{0\mid t}^{(i)}))-\exp(-{\mathcal{E}}_{t}(x_{t}))\right|\leq C\sqrt{\frac{\log(\frac{2}{\delta})}{K}}$ (10)
and
$\left\|\frac{1}{K}\sum_{i}\nabla\exp(-{\mathcal{E}}(x_{0\mid t}^{(i)}))-\nabla\exp(-{\mathcal{E}}_{t}(x_{t}))\right\|\leq C\sqrt{\frac{\log(\frac{2}{\delta})}{K}}$ (11)
@@
  Input: Network $s_{\theta}$, Batch size $b$, Noise schedule $\sigma^{2}_{t}$, Prior $p_{1}$, Num. integration steps $L$, Replay buffer ${\mathcal{B}}$, Max Buffer Size $|{\mathcal{B}}|$ , Num. MC samples $K$.
  while Outer-Loop do
     $\{x_{1}\}^{b}_{i=1}\sim p_{1}(x_{1})$
     $\{x_{0}\}^{b}_{i=1}\leftarrow$ sde_int($(\{x_{1}\}^{b}_{i=1},s_{\theta},L)$Sample}
     ${\mathcal{B}}$ = ({$({\mathcal{B}}\cup\{x_{0}\}^{b}_{i=1})$Update Buffer ${\mathcal{B}}$}
     while Inner-Loop do
        $x_{0}\leftarrow{\mathcal{B}}$.sample() {Uniform sampling from ${\mathcal{B}}$}
        $t\sim\mathcal{U}(0,1)$, $x_{t}\sim\mathcal{N}(x_{0},\sigma^{2}_{t})$
        ${\mathcal{L}}_{\textsc{DEM}}(x_{t},t)=\|{\mathcal{S}}_{K}(x_{t},t)-s_{\theta}(x_{t},t)\|^{2}$
        $\theta\leftarrow\texttt{Update}(\theta,\nabla_{\theta}\mathcal{L}_{\textsc{DEM}}$)
     end while
@@ 3.3 Incorporating symmetries in iDEM
Note that in the first line we used that $x_{0\mid t}^{(i)}\sim\overline{{\mathcal{N}}}(x_{t},\sigma^{2}_{t})$ is equivalent to $g\circ x_{0\mid t}^{(i)}\sim\overline{{\mathcal{N}}}(g\circ x_{t},\sigma^{2}_{t})$. In practice, we can easily implement an equivariant ${\mathcal{S}}_{K}(x_{t},t)$ by replacing the standard normal distribution with a normal distribution that has zero center of mass. Note that a standard normal is already rotation and permutation invariant and an orthogonal action induces no change in volume as the determinant is $1$. Whilst it is not possible to define a translation invariant measure on ${\mathbb{R}}^{d}$ we can still achieve this symmetry by constructing a normal distribution that has zero center of mass. Intuitively, this is a projection of the density in ${\mathbb{R}}^{d}$ to ${\mathbb{R}}^{d-1}$ and this subspace is translation invariant (Köhler et al., 2020; Garcia Satorras et al., 2021; Midgley et al., 2023a). Finally, to use symmetries within our iDEM algorithm we also need the diffusion sampler $s_{\theta}$ to be equivariant to the product group. While this design choice is necessary it does not come with any loss of generality as there always exists an equivariant map between two group invariant distributions on ${\mathbb{R}}^{d}$ (Bose et al., 2021).
@@ 4 Experimental results
We evaluate iDEM on multiple unnormalized densities including synthetic and $\mathrm{SE(3)}\times\mathbb{S}_{n}$-equivariant $n$-body particle systems of varying complexity as examples of scientific applications [^2] Code for iDEM is available at https://github.com/jarridrb/dem.. For our metrics, in Table 2, we report both sample-based metrics, such as 2-Wasserstein that assesses mode coverage, and Effective Sample Size (ESS), as well as the standard negative log-likelihood (NLL). We report additional metrics such as log partition function ($\log{Z}$) and Total Variation (TV) distance in Table 5 in §G.1
@@ 4.1 Main results
Table 2: Sampler performance with mean $\pm$ standard deviation over 3 seeds for negative log-likelihood (NLL), Effective Sample Size (ESS), and 2-Wasserstein metrics ($\mathcal{W}_{2}$). * indicates divergent training. Bold via Welch’s two sample t-test $p<0.1$. See §F.2 for more details.
@@
Algorithm $\downarrow$ NLL ESS ${\mathcal{W}}_{2}$ NLL ESS ${\mathcal{W}}_{2}$ NLL ESS ${\mathcal{W}}_{2}$ NLL ESS ${\mathcal{W}}_{2}$
FAB (Midgley et al., 2023b) 7.14±$\pm 0.01$ 0.653 ±$\pm 0.017$ 12.0±$\pm 5.73$ 7.16±$\pm 0.01$ 0.947 ±$\pm 0.007$ 2.15±$\pm 0.02$ 17.52±$\pm 0.17$ 0.101 ±$\pm 0.059$ 4.35±$\pm 0.01$ 200.32±$\pm 62.3$ 0.063 ±$\pm 0.001$ 18.03±$\pm 1.21$
PIS (Zhang & Chen, 2022) 7.72±$\pm 0.03$ 0.295 ±$\pm 0.018$ 7.64±$\pm 0.92$ 7.19±$\pm 0.01$ 0.901 ±$\pm 0.003$ 2.13±$\pm 0.02$ 47.05±$\pm 12.46$ 0.004 ±$\pm 0.002$ 4.67±$\pm 0.11$
DDS (Vargas et al., 2023) 7.43±$\pm 0.46$ 0.687 ±$\pm 0.208$ 9.31±$\pm 0.82$ 11.27±$\pm 1.24$ 0.408 ±$\pm 0.001$ 2.15±$\pm 0.04$
pDEM (ours) 7.10±$\pm 0.02$ 0.634 ±$\pm 0.084$ 12.20±$\pm 0.14$ 7.44±$\pm 0.05$ 0.547 ±$\pm 0.010$ 2.11±$\pm 0.03$ 18.80±$\pm 0.48$ 0.044 ±$\pm 0.013$ 4.21±$\pm 0.06$
iDEM (ours) 6.96±$\pm 0.07$ 0.734 ±$\pm 0.092$ 7.42±$\pm 3.44$ 7.17±$\pm 0.00$ 0.825 ±$\pm 0.002$ 2.13±$\pm 0.04$ 17.68±$\pm 0.14$ 0.231 ±$\pm 0.005$ 4.26±$\pm 0.03$ 125.86±$\pm 18.03$ 0.106 ±$\pm 0.022$ 16.128±$\pm 0.071$
@@
Table 3: Training time results in hours excluding evaluation time. * denotes divergent training runs.
@@
PIS (Zhang & Chen, 2022) 4.11 11.29 17.36 *
DDS (Vargas et al., 2023) 1.81 5.65 * *
pDEM (ours) 0.36 1.40 1.79 *
@@ 6 Conclusion
J.R., N.M., and A.T. initially conceived the idea of a stochastic off-policy continuous regression objective. A.J.B., T.A., G.G., and A.T. independently approached the problem through the lens of active inference with a generative model in a bi-level iterative scheme. Experiments were primarily led by T.A., J.R., A.T. (iDEM) and J.R,S.M., P.L., C.H.L., M.S. (ablations and baselines). G.G., A.J.B., and N.M. led the development of the theory. S.R., G.G., and Y.B. guided the project. A.J.B. and N.M. drove the writing of the paper, with contributions from all other authors. All authors contributed to designing the experiments.
@@ D.5 Boltzmann generators
A Boltzmann generator (BG) samples from $\mu_{\text{target}}$ by combining an exact-likelihood model $q_{\theta}$, typically a Normalizing Flow, and an algorithm to reweight model generated samples using $\mu_{\text{target}}$. During training, a BG’s flow is learned by minimizing a convex combination of both the forward and reverse KL-divergence. Thus, to compute the full training loss we need a dataset of ground truth samples from $\mu_{\text{target}}$ as it is required under the forward KL. As a result, BG’s are ill-suited in settings with limited or no data as training only with the reverse KL is prone to mode-seeking behavior. Given a trained flow $q_{\theta}$ we can reweight observables $f(x)$ to be under $\mu_{\text{target}}$ by using the IS estimator.
@@ F.1 Experimental Setup
F.2 Metrics reported in Table 2 and Table 5
Tab. 2 depicts the main experimental sample quality results for various methods on all datasets and additional metrics are reported in Tab. 5. In this section, we explain the methodology for each experiment. We note that it was difficult to train PIS and DDS for the high-dimensional Lennard-Jones tasks. 2/3 seeds of DDS diverged for LJ-13.
@@ F.2 Metrics reported in Table 2 and Table 5
2-Wasserstein distance $\mathcal{W}_{2}$. We also use the standard 2-Wasserstein distance between empirical samples from the sampler and the ground truth dataset. The 2-Wasserstein distance is defined as
@@
where $\pi$ is the transport plan with marginals constrained to $\mu$ and $\nu$ respectively. In practice, we use the Hungarian algorithm as implemented in the Python optimal transport package (POT) (Flamary et al., 2021) to solve this optimization for discrete samples. For simplicity, we use the Euclidean ground distance $d(x,y)=\|x-y\|_{2}$.
Effective Sample Size (ESS). To measure ESS we first evaluate log importance weights given by using our model $p_{model}(x)$ as a proposal. We estimate the ESS by
$ESS=\frac{n}{\sum_{i=1}^{n}w_{i}^{2}}$
where $w_{i}=\frac{\exp(-\mathcal{E}(x_{i}))/p_{model}(x_{i})}{\sum_{j=1}^{n}\exp(-\mathcal{E}(x_{j}))/p_{model}(x_{j})}$, which is equivalent to the Softmax of the log-probability ratios. We note that this is sometimes known as the normalized ESS, as we normalize the effective sample size to the fraction of the test set size.
Total Variation. The total variation metric on the Gaussian mixture model dataset is taken over 200 bins in each dimension (2002) total bins. This is possible in low dimensions but does not scale well to high dimensions, requiring an exponential number of bins. Therefore for the larger equivariant datasets, we take the total variation distance over the distribution of the interatomic distances of the particles, see e.g., Fig. 4, which scales well with dimension and shows how well the distribution of energies matches that of the test data.
Log partition function ( $\log{Z}$ ). To compute $\log Z$ we use an importance sampling estimate with the proposal density, $q(x)$, given by the OT-CFM model as
$\log Z$ $=\log\mathbb{E}_{x\sim q(x)}\left[\frac{\exp(-\mathcal{E}(x))}{q(x)}\right]$ (33)
$\geq\mathbb{E}_{x\sim q(x)}[-\mathcal{E}(x)-\log q(x)]$ (34)
which yields a lower bound on the true log partition function. Note that for GMM we know the true value of $\log{Z}$, however, this isn’t the case for the equivariant tasks. For those tasks, to compare the values of $\log{Z}$ from the different methods, we use the fact that our estimate is a lower bound of the true value and favour the method with the largest estimate.
@@ F.4.1 Gaussian Mixture Model
We use a test set of size 1000 sampled with torch.random.seed(0) following prior work.
@@ G.1 Additional metrics
We report here additional quantitative metrics to supplement the main results presented in §4.1. In Tab. 5, we report the Total Variation (TV) as well as the log partition function ($\log{Z}$) values.
Table 5: Sampler performance with mean $\pm$ standard deviation over 3 seeds for Total Variation (TV), and log partition function ($\log Z$). * indicates divergent training and entries with refer to settings where only 1 of 3 runs converged. For TV, we bold via Welch’s two-sample t-test $p<0.1$. For $\log{Z}$, we bold the method with the largest value, as our estimate is a lower bound on the true $\log{Z}$.
Energy $\rightarrow$ GMM ($d=2$) DW-4 ($d=8$) LJ-13 ($d=39$) LJ-55 ($d=165$)
Algorithm $\downarrow$ TV $\log{Z}$ TV $\log{Z}$ TV $\log{Z}$ TV $\log{Z}$
FAB (Midgley et al., 2023b) 0.88±$\pm 0.02$1.165 ±$\pm 0.164$ 0.09±$\pm 0.00$ 29.602 ±$\pm 0.019$ 0.04±$\pm 0.00$ 4.35±$\pm 0.01$ 0.24±$\pm 0.09$ 32.809
PIS (Zhang & Chen, 2022) 0.92±$\pm 0.01$2.243 ±$\pm 0.070$ 0.09±$\pm 0.00$ 29.599 ±$\pm 0.009$ 0.25±$\pm 0.01$ 46.685 ±$\pm 1.471$
DDS (Vargas et al., 2023) 0.82±$\pm 0.02$0.358 ±$\pm 0.209$ 0.16±$\pm 0.01$ 28.382 ±$\pm 0.158$
pDEM (ours) 0.82±$\pm 0.02$0.370 ±$\pm 0.005$ 0.13±$\pm 0.00$ 29.191 ±$\pm 0.036$ 0.06±$\pm 0.02$ 32.450 ±3plus-or-minus3$\pm$ 3± 3.191 * *
iDEM (ours) 0.82±$\pm 0.01$0.340 ±$\pm 0.075$ $\pm 0.01$ 29.567 ±$\pm 0.014$ 0.04±$\pm 0.01$ 49.969 ±$\pm 2.784$ 0.09±$\pm 0.01$ 273.167 ±$\pm 22.226$
In Tab. 6 we report the NLL values that are computed between the original method and the CFM model that is trained using the samples of each method.
@@ G.3 Further ablations
$=-\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{x_{0|t}\sim\mathcal{N}(x_{t},\sigma_{t}^{2})}\left[\mathcal{E}(x_{0|t})(x_{0|t}-x_{t})\right],$
@@
We investigate the behaviour of each estimate in Figure 10, examining the MSE between the score estimated by each technique and the true score on the GMM task as a function of the number of MC samples. Unfortunately, we observe that despite the Jensen estimate’s attractiveness the score estimates it yields have a large MSE which does not readily reduce as the number of MC samples increases indicating a significant bias. On the other hand, the unbiased estimate suffers from extremely large variance due to it having to work directly with the exponentiated energies $\exp(-\mathcal{E}(x_{t}))$. This results in every estimate before 500 MC samples resulting in exclusively NaN or infinite estimates, rendering the estimate unusable in practice. The only estimate which is well-behaved and approaches the true score is the LogSumExp estimate, whose use we advocate throughout this work.
@@
In Fig. 11 (left) we study the log-bias as a function of ${\mathcal{E}}$, by taking a point from a GMM mode and linearly following a direction away from the data itself which corresponds to $\exp(-{\mathcal{E}}(x))\approx 0$. As observed, we notice we need a few MC samples in high-data regions. Additionally, the log-bias increases further from data for a fixed $K$ and drops when we increase $K$. This result empirically confirms §3.1 where we need more MC samples to have a proper estimation of the score.
@@
Figure 12: Plot of the Left: bias-squared Middle: mean squared error (MSE) and Right: cosine similarity of the mean estimated score to the ground truth score as a function of the number of dimensions of the 40-GMM. These are plotted for $x_{t}$ sampled at three noise levels $\sigma(t)\in\{2,4,8\}$. The standard deviations for all the values are over $10$ seeds with the variance estimated over $128$ samples.
Estimator quality as a function of dimension. We examine the quality of the estimator as a function of dimensions, with a closed-form score for the $40$-GMM task. Specifically, we analyze the bias and the MSE (capturing variance) of the estimator as a function of increasing number of dimensions, across various noise levels, $\sigma(t)$ in Fig. 12 (left, middle). As expected, we find that the bias and MSE increase as a function of dimensions. In line with our expectations and previous findings shown in Fig. 5 (right), we note that the bias and MSE are both higher at higher noise levels. In Fig. 12 (right), we measure the cosine similarity of the mean estimated score to the true score. As this value is very close to $1$ for the dimensions we analyze, we conclude that, interestingly, the estimator error is almost exclusively in the magnitude and not in the direction. This interesting finding is also aligned with our experiments, where we observed that we were able to get good samples, even in high-dimensional tasks, simply by clipping the estimated score. Overall, we hypothesize that overestimating the magnitude of the score early in the diffusion process (i.e., at large values of $\sigma(t)$) does not pose a significant problem as long as the estimated scores still point towards the data density and that later in the diffusion process (i.e., at smaller values of $\sigma(t)$) we are able to get accurate estimates. The reason is that although using the overestimated scores we might move towards the data density too quickly or even overshoot it, by taking small enough steps we will eventually enter a region where our estimates are accurate and we can converge towards the data density.
Finally, to demonstrate that the errors in the magnitude of the estimated score are manageable even in a very high-dimensional setting, we scale up the $40$-GMM task from $2$ to $10000$ dimensions. As we want to examine the quality of the estimator itself, we use the estimator directly, instead of a network, to generate samples. Although clipping is necessary to achieve good samples, the estimator is able to achieve good performance over a wide range of clipping values, from $10^{4}$ to $10^{8}$. In Fig. 13, we demonstrate this by plotting the true and predicted distributions on the first two principal components using the estimator, with norm clipped to $10^{4}$ and $10^{8}$ respectively.
Figure 13: Generated and true samples for the 40-GMM task in 10,000 dimensions plotted over the first two principal components. The samples are generated using the score estimator with norm clipped to Left: $10^{4}$ and Right: $10^{8}$.